# Generating Shrinkers by Mean Curvature Flow

### v2

@@ 7. Shrinkers with one end
    $\Sigma_{g}\setminus\{z=0\}$ has genus 00.
@@
Existence of shrinkers with properties (1)–(4) was proved (by surgery), for all sufficiently large $g$, by Kapouleas, Kleene, and Moeller [kkm], and, independently, by X. H. Nguyen [nguyen]. It was subsquently proved for all $g$ by Buzano, Huy The Nguyen, and Schulz using minimax [buzano]. However, none of those works proved that any of the shrinkers arose as blowups of the mean curvature flow of a compact, smoothly embedded initial surface.
Figure 6. The piecewise smooth surface
$\Delta_{R}\cup(\{z>0\}\cap\partial B(0,R))$
in the case $g=1$. Its boundary consists of a pair of diameters, and a pair of circular arcs.
@@
For $20\leq R<$ 2, let
$\Delta_{R}$ =$=\{(r\cos\theta,r\sin\theta,0):\text{$0\leq r\leq R$ and $\sin((g+1)\theta)\geq 0$}\}$
@@
Note that $S$ can be obtained by taking the union of the disk $D$ and the sphere $\partial B(0,R)$ and then doing surgery along the circle of intersection. The Euler characteristic of the disjoint union is $1+2=3$. The surgery does not change the number of edges (in a suitable triangulation); it merely changes the way that faces are attached to edges. But the surgery does reduce the number of vertices by $2(g+1)$, since $2(g+1)$ vertices on the equator of the sphere get identified with the corresponding points in $D$. Thus
@@
Now let $M$ be any surface in $\mathscr{M}_{g}^{\prime}$. We orient $M$ so that the unit normal at $(2,0,0)$ is $\mathbf{e}_{3}$. It follows that the unit normal at the origin is $-\mathbf{e}_{3}$:
@@
By Theorems 33 and 34, the flow is regular at all points of $\tilde{Q}$ for all times $t\in[0,T_{\textnormal{pos}})$. In particular, the points $(0,0,0)$ and $(2,0,0)$ are such points. At each of those two points, two or more great circles in $\widetilde{Q}$ intersect transversely. Since $\widetilde{Q}\subset M(t)$ for all $t\in[0,T_{\textnormal{pos}})$, we see that the tangent plane to $M(t)$ at each of those two points is horizontal. Thus since (21) holds for $t=0$ (by (20)), it holds for all $t\in[0,T_{\textnormal{pos}})$. ∎
@@
By Theorem 37 below, we can choose $\Sigma_{g}$ to have entropy $<\delta$. By Theorem 38, given that entropy bound, $\Sigma_{g}$ has the asserted behavior as $g\to\infty$. ∎
@@
Note that there exist $M\in\mathscr{M}^{\prime}_{g}$ that are arbitrarily close to $\Psi_{r}$ in the weak sense. That is, there exists a sequence $M_{i}\in\mathscr{M}^{\prime}_{g}$ such that $M_{i}$ converges weakly to $\Psi_{r}$.
@@
    $M_{i}$ converges weakly with with multiplicity $1$ to $\{z=0\}$
@@
is a mean curvature flow in $\mathbf{R}^{2}\cup Z_{\infty}$, but with the dilated metric $\gamma_{\lambda_{i}}$, where $\lambda_{i}:=(T_{i})^{-1/2}$.
@@
where $\epsilon_{i}\to 0$. (The $\epsilon_{i}$ is there because $M_{i}^{\prime}(\cdot)$ is mean curvature flow for a non-flat metric; and $\epsilon_{i}\to 0$ because the metric is converging to smoothly to the flat metric.)
@@ 8. Shrinkers with three ends
    $\Sigma_{g}\setminus\{z=0\}$ has genus 00.
@@
Figure 7. Conjectured qualitative appearance of the shrinker $\Sigma_{g}$ in case $g=1$. The surface shown has genus $2=2g$ and it seems to have $3$ ends. However, we know that $\Sigma_{g}$ has genus $2g$ and $3$ ends only when $g$ is large. Figure courtesy of Matthias Weber. (This figure is not based on a numerical approximation of a shrinker.)
@@
    In Theorem 55 and Proposition 58, we show, for all sufficiently large $g$, that $\Sigma_{g}$ has the remaining properties asserted in Theorem 39: it has genus $2g$, it has exactly $3$ ends, and it converges as $g\to\infty$ to the plane $\{z=0\}$ with multiplicity $3$.
@@
$W_{k}:=\frac{\pi k}{g+1}<\theta<\frac{\pi(k+1)}{g+1}.$
Let $\Omega(s)$, $10\leq s<$ 1, be a smooth, $1$-parameter family of nested, closed convex regions in $W_{0}\cap D(R)$ such that $\Omega(0)$ is a single point and such that $\cup_{s}\Omega(s)$ is all of $W_{0}\cap D(R)$.
@@
For $10<s<$ 1, we do surgery on $D\cup A^{+}_{s}\cup A^{-}_{s}$ as follows. For each $W_{k}$ with $k$ even, we connect $D$ to $A^{+}_{s}$ be a neck in the region $W_{k}$. Likewise, for each $W_{k}$ with $k$ odd, we connected $D$ to $A^{-}_{s}$ by a neck in $W_{k}$. We do this in a way that preserves the $G_{g}$ symmetry, so it suffices to describe the surgery in $W_{0}$. The surgery in $W_{0}$ is the following: remove $\Omega(s)$ from $D$, remove the corresponding region
@@
where the second-to-last inequality follows from (23). The same bound holds for $\operatorname{area}_{\,\gamma}(\operatorname{neck}[s])$ since $\gamma_{ij}\leq\delta_{ij}$ everywhere. Consequently, since $M[s]$ has $4(g+1)$ necks,
@@
We fix an $R$ for which this last expression is $<5\pi$. Thus
@@
From the construction, it is straightforward to check that $M^{\prime}[s]$ has genus $4g$ and that $M^{\prime}[s]\setminus\{z=0\}$ has genus 00. Thus $M^{\prime}[s]\in\mathscr{M}_{g}^{\prime\prime}$. ∎
@@
=$=\begin{cases}3\pi r^{2}&\text{if $r<R$}\\ \pi r^{2}+6\pi R^{2}&\text{if $r>R$}.\end{cases}$
@@
Let $\mathscr{F}=\mathscr{F}_{g}$ be the family of smooth surfaces $M^{\prime}[s]$ constructed in the proof of Proposition 40, where the smoothing has been done in accord with Remarks 41 and 43.
@@
$\nu_{M(t)}(0,0,0)=\begin{cases}\mathbf{e}_{3}&\text{if $n(t)$ is even,}\\ -\mathbf{e}_{3}&\text{if $n(t)$ is odd.}\end{cases}$
@@
by Lemma 22. Thus $n(t)$ is even for all regular times $t\in[0,T_{\textnormal{pos}})$. ∎
@@
By the Initial Regularity Theorem (Theorem 8), there is a time $\tau>0$ such that for all sufficiently large $n$, the flow $M_{n}(\cdot)$ has no singularities on $Z$ during the time interval $[0,\tau]$. In particular, for $t\in[0,\tau]$, $M_{n}(t)\cap Z$ consists of exactly $5$ points, and those points depend smoothly on $t$. Thus, by replacing $\tau$ by a smaller $\tau>0$, we can assume that
@@
    $\Sigma\setminus\{z=0\}$ has genus 00,
@@
Trivially, $\Sigma_{g}$ is $G_{g}$-invariant. Note that $M(t)\setminus\{z=0\}$ has genus 00 for all regular $t<T_{\textnormal{pos}}$ by Theorem 45. Hence $\Sigma\setminus\{z=0\}$ has genus 00. By Theorem 35, $\Sigma\cap\{z=0\}=Q_{g}$ and $\Sigma$ has genus $g$ or $2g$. Finally, the uniform entropy bound follows from Remark 43 and Huisken monotonicity. ∎
@@
that admits a strong orientation $\nu$ with the following property: for each $t$, if the origin is a regular point at time $t$, then $\nu(O,t)=-\mathbf{e}_{3}$.
@@
By the Local Regularity Theorem (Theorem 7), $\tilde{M}(\cdot)$ is special since each $M_{n}(\cdot)$ is.
@@
occurs as a tangent flow to $M(\cdot)$ at the spacetime point $(0,T)$, so it also occurs as a tangent flow to $\tilde{M}(\cdot)$ at that point, since the two flows coincide up to time $T$. Thus there is a sequence $\lambda_{n}\to\infty$ such that the parabolically dilated flows
@@
By passing to a subsequence (if necessary) we can assume that the flows converge for all $t$ a flow
@@
Since the $\tilde{M}(\cdot)$ is special, so are the flows (35), and therefore (by the Local Regularity Theorem, Theorem 7) so is the flow (36). ∎
@@
$M^{\prime}(g)=|t|^{-1/2}\Sigma\quad(t<0).$
Note that $M^{\prime}(0)$, the limit of $|t|^{-1/2}\Sigma$ as $t\to 0$ is the $P$, the plane $\{z=0\}$ with multipicity $1$. Hence $M^{\prime}(t)=P$ for all $t\geq 0$. That is, $M^{\prime}(\cdot)$ is the flow
(37) $t\in\mathbf{R}\mapsto\begin{cases}|t|^{-1/2}\Sigma&(t<0),\\ \Sigma&(t\geq 0).\end{cases}$
@@
Note that at each time $t=(-1)$, the singular set is the circle $J:=S\cap P$. Thus, by specialness, there is a closed set $K$ such that $K\setminus J$ is a smooth manifold-with-boundary. At the origin, $-\mathbf{e}_{3}$ points out of $K$. It follows that $K$ consists of the closure of the region above $P$ and inside $S$, together with the closure of region the region below $P$ and outside $S$. See Fig. 9.
Figure 9. The region $K$ (in the case that $S$ is a sphere)
By specialness, at time $t=1$,
Thus $\nu(p,-1)=\mathbf{e}_{3}$ at all points $p$ in $\{z=0\}$ that lie outside $S$.
@@
$\nu(p,t)=\mathbf{e}_{3}$
for all $p$ in $\{z=0\}$ and $t>0$. In particular, $\nu(0,t)=\mathbf{e}_{3}$ for all $t>0$, contrary to specialnesss of $M^{\prime}(\cdot)$. ∎
@@ 9. Limits of Shrinkers
The proof of [brendle]*Theorem 2 shows that $\Sigma_{g}^{+}$ is connected. Since $\Sigma_{g}^{+}$ has genus 00, $\Gamma_{g}$ divides $\Sigma_{g}^{+}$ into two connected components $U_{g}$ and $U_{g}^{\prime}$. Note that $U_{g}$ and $U_{g}^{\prime}$ are each $G_{g}^{+}$-invariant. Since $\Sigma_{g}$ is connected, the closure of at least one component, which we may assume is $U_{g}^{\prime}$, contains points in $Q_{g}$. It follows that $\overline{U_{g}^{\prime}}$ contains an entire ray of $Q$. By $G_{g}^{+}$ invariance, $\overline{U_{g}^{\prime}}$ contains all of $Q$. It follows that $\Gamma_{g}$ divides $\Sigma_{g}$ into two components, namely $U_{g}$ and
@@
By passing to a subsequence, we can assume that the $\Sigma_{g}$ converge as sets to a limit set $\Sigma$. We can also assume that there is a closed set $K$ (the curvature blowup set) such the convergence $\Sigma_{g}$ to $\Sigma$ is smooth (perhaps with multiplicity) in compact subsets of $\mathbf{R}^{3}\setminus K$ and such that
@@
(38) $\Sigma$ =$=\{z=0\}\cup\Sigma^{+}\cup\Sigma^{-}$
=$=\{z=0\}\cup\Sigma^{+}\cup\rho_{X}\Sigma^{+},$
@@
By assumption, $\Sigma_{g}^{+}:=\Sigma_{g}\cap\{z>0\}$ has genus 00. By the compactness theorem [white-compact]*Theorem 1.1 for minimal surfaces of locally bounded area and genus, $\Sigma^{+}$ is smooth and $K\cap\{z>0\}$ is a discrete subset of $\{z>0\}$. The proof of [brendle]*Theorem 2 shows that $\Sigma^{+}$ is unstable (as a minimal surface for the shrinker metric) and connected. Thus (again by the compactness theorem [white-compact]*Theorem 1.1), the convergence of $\Sigma_{g}^{+}$ to $\Sigma^{+}$ is smooth and with multiplicity $1$, and therefore $K\cap\{z>0\}$ is empty. Therefore, by symmetry,
@@
Suppose $\Sigma^{+}$ is nonempty. let $C^{\prime}$ be a horizontal circle in $\Sigma^{+}$. Let $C_{q}^{\prime}$ be smooth, simple closed curves in $\Sigma_{q}$ that converge smoothly to $C^{\prime}$. By Lemma 56, $C_{q}^{\prime}$ bounds a region $E_{n}$ in $\Sigma_{n}$ that is either a disk or an annulus, that lies in $\{z>0\}$, and that is properly embedded in $\mathbf{R}^{3}$. Passing to the limit, we see that $C^{\prime}$ bounds a region $E$ in $\Sigma^{+}$ that is either a disk or an annulus, and that is properly embedded in $\mathbf{R}^{3}$.
The other end $E^{\prime}$ of $\Sigma^{+}$ must diverge to infinity, converge to the origin, or converge to a circle $C(r)$ in the plane $\{z=0\}$. The end $E^{\prime}$ cannot diverge to infinity, because then $\Sigma^{+}$ would be shrinker (without boundary) in $\{z>0\}$. The end $E^{\prime}$ cannot converge the origin, for if it did, then $\Sigma^{+}\cup\{0\}$ and $\{z=0\}$ would violate the strong maximum principle at the origin. (One can invoke various versions of the maximum principle here. For example, the classical strong maximum principle suffices, because $\Sigma^{+}\cup\{0\}$ is a smooth minimal surface with respect to the shrinker metric by [structure]*Corollary 2.7.)
Thus the end $E^{\prime}$ converges to a circle $C(r)$ in the plane $\{z=0\}$. This completes the proof of Claim 1. ∎
@@
By rotational symmetry by $\pi$ about $X$, If $L$ is a closed vertical line segment centered at $p$, then $L\cap\Sigma_{g}$ has an odd number of points (namely, the point $p$ and equal numbers of points above and below $p$.) ∎
@@
$-m\mathbf{e}_{1}+m^{\prime}\mathbf{e}_{1}+\nu^{+}+\nu^{-}=0.$
@@
If $m=1$, the convergence is smooth by Allard’s Regularity Theorem. Thus we assume that $m>1$.
For all sufficiently large $g$, $\Sigma_{g}\cap\partial B(0,r)$ consists of $m$ simple closed curves. Let $\Gamma_{g}^{0}$ be the curve that passes through the points of $Q_{g}\cap\partial B(0,r)$. Let $\Gamma_{g}$ be one of the other components. By symmetry, we may assume that $\Gamma_{g}$ lies in $\{z>0\}$. Let $\Sigma_{g}$ be the component of $\Sigma_{g}\setminus\Gamma_{g}$ that lies in $\{z>0\}$. (The component exists by Lemma 56). Since $\Sigma^{g}\setminus B(0,r)$ is connected and contains points in $\{z\leq 0\}$, we see that $\Sigma_{g}$ is contained in $B(0,r)$.

### v3

@@ 7. Shrinkers with one end
    $\Sigma_{g}\setminus\{z=0\}$ has genus 0.
@@
Existence of shrinkers with properties (1)–(4) was proved (by surgery), for all sufficiently large $g$, by Kapouleas, Kleene, and Moeller [kkm], and, independently, by X. H. Nguyen [nguyen]. It was subsquently proved for all $g$ by Buzano, Huy The Nguyen, and Schulz using minimax [buzano]. Those papers also proved that, as $g\to\infty$, the shrinkers have the behavior stated in Theorem 40. However, none of those papers proved that any of the shrinkers arise as blowups of the mean curvature flow of a compact, smoothly embedded initial surface.
Figure 6. The piecewise smooth surface Δ_R∪({z¿0}∩∂B(0,R)) in the case $g=1$. Its boundary consists of a pair of diameters, and a pair of circular arcs.
@@
For $0\leq R<2$, let
$\Delta_{R}$ $=\{(r\cos\theta,r\sin\theta,0):\text{$0\leq r\leq R$ and $\sin((g+1)\theta)\geq 0$}\}$
@@
Note that $S$ can be obtained by taking the union of the disk $D$ of Euclidean radius 2 in the plane $\{z=0\}$ and the sphere $\partial B(0,R)$ and then doing surgery along the circle of intersection. The Euler characteristic of the disjoint union is $1+2=3$. The surgery does not change the number of edges (in a suitable triangulation); it merely changes the way that faces are attached to edges. But the surgery does reduce the number of vertices by $2(g+1)$, since $2(g+1)$ vertices on the equator of the sphere get identified with the corresponding points in $D$. Thus
@@
Now let $M$ be any surface in $\mathscr{M}_{g}^{\prime}$. We orient $M$ so that the unit normal at $(2,0,0)$ is $\mathbf{e}_{3}$. It follows that the unit normal at the origin is $\mathbf{e}_{3}$:
@@
By Theorems 37 and 38, the flow is regular at all points of $\tilde{Q}$ for all times $t\in[0,T_{\textnormal{pos}})$. In particular, the points $(0,0,0)$ and $(2,0,0)$ are such points. At each of those two points, two or more great circles in $\widetilde{Q}$ intersect transversely. Since $\widetilde{Q}\subset M(t)$ for all $t\in[0,T_{\textnormal{pos}})$, we see that the tangent plane to $M(t)$ at each of those two points is horizontal. Thus since (27) holds for $t=0$ (by (26)), it holds for all $t\in[0,T_{\textnormal{pos}})$. ∎
@@
By Theorem 41 below, we can choose $\Sigma_{g}$ to have entropy $\delta$. By Theorem 42, given that entropy bound, $\Sigma_{g}$ has the behavior asserted in the statement of Theorem 40, as $g\to\infty$. ∎
@@
Note that there exist $M\in\mathscr{M}^{\prime}_{g}$ that are arbitrarily close to $\Psi(r)$ in the weak sense. That is, there exists a sequence $M_{i}\in\mathscr{M}^{\prime}_{g}$ such that $M_{i}$ converges weakly to $\Psi(r)$.
@@
    $M_{i}$ converges weakly with multiplicity $1$ to $\{z=0\}$
@@
We claim that
(29) $T_{i}\to 0.$
For suppose not. Then (after passing to a subsequence) $T_{i}$ converges to a limit $T\in(0,\infty]$. Since $M_{i}$ converges weakly with multiplicity $1$ to $\{z=0\}$, the flow $t\in[0,T_{i}]\mapsto M_{i}(t)$ converges weakly to the multiplicity-one flow
$t\in[0,T]\mapsto\{z=0\}.$
By local regularity theory [white-local], the convergence is smooth on $\epsilon,T]$ for every $\epsilon>0$. In particular, $M_{i}(\epsilon)$ is a smooth graph for all sufficiently large $i$. But that contradicts Claim 1 in the proof of Theorem 40, thus proving (29).
@@
is a mean curvature flow in $\mathbf{R}^{3}\cup Z_{\infty}$, but with the dilated metric $\gamma_{\lambda_{i}}$, where $\lambda_{i}:=(T_{i})^{-1/2}$.
@@
where $\epsilon_{i}\to 0$. (The $\epsilon_{i}$ is there because $M_{i}^{\prime}(\cdot)$ is mean curvature flow for a non-flat metric; and $\epsilon_{i}\to 0$ because the metric is converging to smoothly to the flat metric by (29).)
@@ 8. Shrinkers with three ends
    $\Sigma_{g}\setminus\{z=0\}$ has genus 0.
@@
Figure 7. Conjectured qualitative appearance of the shrinker $\Sigma_{g}$ in case $g=1$. The surface shown has genus $2=2g$ and it seems to have 33 ends. However, we know that $\Sigma_{g}$ has genus $2g$ and 33 ends only when $g$ is large. Figure courtesy of Matthias Weber. (This figure is not based on a numerical approximation of a shrinker.)
@@
    In Theorem 59 and Proposition 62, we show, for all sufficiently large $g$, that $\Sigma_{g}$ has the remaining properties asserted in Theorem 43: it has genus $2g$, it has exactly 33 ends, and it converges as $g\to\infty$ to the plane $\{z=0\}$ with multiplicity $3$.
@@
$W_{k}:=\frac{\pi k}{g+1}<\theta<\frac{\pi(k+1)}{g+1},$
where $0\leq k\leq 2g+1$. Let $\Omega(s)$, $0\leq s<1$, be a smooth, $1$-parameter family of nested, closed convex regions in $W_{0}\cap D(R)$ such that $\Omega(0)$ is a single point, such that $\cup_{s}\Omega(s)$ is all of $W_{0}\cap D(R)$, and such that each $\Omega(s)$ is symmetric under reflection in the plane $P_{\pi/2(g+1)}$.
@@
For $0<s<1$, we do surgery on $D\cup A^{+}_{s}\cup A^{-}_{s}$ as follows. For each $W_{k}$ with $k$ even, we connect $D$ to $A^{+}_{s}$ by a neck in the region $W_{k}$. Likewise, for each $W_{k}$ with $k$ odd, we connected $D$ to $A^{-}_{s}$ by a neck in $W_{k}$. We do this in a way that preserves the $G_{g}$ symmetry, so it suffices to describe the surgery in $W_{0}$. The surgery in $W_{0}$ is the following: remove $\Omega(s)$ from $D$, remove the corresponding region
@@
where the second-to-last inequality follows from (30). The same bound holds for $\operatorname{area}_{\,\gamma}(\operatorname{neck}[s])$ since $\gamma_{ij}\leq\delta_{ij}$ everywhere (14).
Consequently, since $M[s]$ has $4(g+1)$ necks,
@@
We fix an $R$ for which this last expression is $5\pi$. Thus
@@
From the construction, it is straightforward to check that $M^{\prime}[s]$ has genus $4g$ and that $M^{\prime}[s]\setminus\{z=0\}$ has genus 0. Thus $M^{\prime}[s]\in\mathscr{M}_{g}^{\prime\prime}$. ∎
@@
$=\begin{cases}3\pi r^{2}&\text{if $r<R$}\\ \pi r^{2}+6\pi R^{2}&\text{if $r>R$}.\end{cases}$
@@
Let $\mathscr{F}=\mathscr{F}_{g}$ be the family of smooth surfaces $M^{\prime}[s]$ constructed in the proof of Proposition 44, where the smoothing has been done in accord with Remarks 45 and 47.
@@
(37) $\nu_{M(t)}(0,0,0)=\begin{cases}\mathbf{e}_{3}&\text{if $n(t)$ is even,}\\ -\mathbf{e}_{3}&\text{if $n(t)$ is odd.}\end{cases}$
@@
by Lemma 25. Thus, for all regular times $t\in[0,T_{\textnormal{pos}})$, $n(t)$ is even and so $\nu_{M(t)}(0,0,0)=\mathbf{e}_{3}$ by (37). ∎
@@
By the Initial Regularity Theorem (Theorem 9), there is a time $\tau>0$ such that for all sufficiently large $n$, the flow $M_{n}(\cdot)$ has no singularities on $Z$ during the time interval $0,\tau]$. In particular, for $t\in[0,\tau]$, $M_{n}(t)\cap Z$ consists of exactly $5$ points, and those points depend smoothly on $t$. Thus, by replacing $\tau$ by a smaller $\tau>0$, we can assume that
@@
    $\Sigma\setminus\{z=0\}$ has genus 0,
@@
Trivially, $\Sigma_{g}$ is $G_{g}$-invariant. Note that $M(t)\setminus\{z=0\}$ has genus 0 for all regular $t<T_{\textnormal{pos}}$ by Theorem 49. Hence $\Sigma\setminus\{z=0\}$ has genus 0. By Theorem 39, $\Sigma\cap\{z=0\}=Q_{g}$ and $\Sigma$ has genus $g$ or $2g$. Finally, the uniform entropy bound follows from Remark 47 and Huisken monotonicity. ∎
@@
that admits a strong orientation $\nu$ with the following property: for each $t$, if the origin is a regular point at time $t$, then $\nu(O,t)=\mathbf{e}_{3}$.
@@
[…]
@@
[…]
@@
By passing to a subsequence (if necessary) we can assume that the flows (43) converge for all $t$ to a flow
@@
[…]
@@
$M^{\prime}(g)=|t|^{-1/2}\Sigma\quad(t\leq 0).$
As a set, $M^{\prime}(0)$ is $P$ if $S$ is the sphere and $P\cup Z$ if $S$ is the cylinder. In either case, the corresponding Radon measure is the plane $P$ with multiplicity one. It follows (by unit regularity) that $M^{\prime}(t)$ is the plane $P$ with multiplicity one for all $t>0$.
That is, $M^{\prime}(\cdot)$ is the flow
(45) $t\in\mathbf{R}\mapsto\begin{cases}|t|^{-1/2}\Sigma&(t\leq 0),\\ P&(t>0).\end{cases}$
@@
Note that at time t=1t=-1, the singular set of (45) is the circle $J:=S\cap P$. Thus, by specialness, there is a closed set $K$ such that $K\setminus J$ is a smooth manifold-with-boundary. At the origin, $\mathbf{e}_{3}$ points out of $K$. It follows that $K$ consists of the closure of the region below $P$ and inside $S$, together with the closure of region the region above $P$ and outside $S$. See Fig. 9.
Figure 9. The region $K$ (in the case that $S$ is a sphere).
Thus $\nu(p,-1)=-\mathbf{e}_{3}$ at all points $p$ in $\{z=0\}$ that lie outside $S$.
@@
$\nu(p,t)=-\mathbf{e}_{3}$
for all $p$ in $\{z=0\}$ and $t>0$. In particular, $\nu(0,t)=-\mathbf{e}_{3}$ for all $t>0$, contrary to specialnesss of $M^{\prime}(\cdot)$. ∎
@@ 9. Limits of Shrinkers
The proof of [brendle]*Theorem 2 shows that $\Sigma_{g}^{+}$ is connected. Since $\Sigma_{g}^{+}$ has genus 0, $\Gamma_{g}$ divides $\Sigma_{g}^{+}$ into two connected components $U_{g}$ and $U_{g}^{\prime}$. Now $Q=Q_{g}$ is in the closure of $\Sigma_{g}\setminus Q$, and, therefore, by symmetry, in the closure of $\Sigma_{g}^{+}$. Thus the closure of one of the components, say $\overline{U_{g}^{\prime}}$, of $\Sigma_{g}\setminus\Gamma_{g}$, contains a point of $Q\setminus\{O\}$. Hence $\overline{U_{g}^{\prime}}$ contains the corresponding connected component of $Q_{g}\setminus\{O\}$. By $G_{g}^{+}$ symmetry, $\overline{U_{g}^{\prime}}$ contains all of $Q_{g}\setminus\{O\}$ and therefore all of $Q_{g}$. It follows that $\Gamma_{g}$ divides $\Sigma_{g}$ into two components, namely $U_{g}$ and
@@
By passing to a subsequence, we can assume that the $\Sigma_{g}$ converge as sets to a limit set $\Sigma$ that is rotationally invariant about $Z$. We can also assume that there is a closed set $K$ (the curvature blowup set) such the convergence $\Sigma_{g}$ to $\Sigma$ is smooth (perhaps with multiplicity) in compact subsets of $\mathbf{R}^{3}\setminus K$ and such that
@@
(46) $\Sigma$ $=\{z=0\}\cup\Sigma^{+}\cup\Sigma^{-}$
$=\{z=0\}\cup\Sigma^{+}\cup\rho_{X}\Sigma^{+},$
@@
By assumption, $\Sigma_{g}^{+}:=\Sigma_{g}\cap\{z>0\}$ has genus 0. By the compactness theorem [white-compact]*Theorem 1.1 for minimal surfaces of locally bounded area and genus, $\Sigma^{+}$ is smooth and $K\cap\{z>0\}$ is a discrete subset of $\{z>0\}$. The proof of [brendle]*Theorem 2 shows that $\Sigma^{+}$ is unstable (as a minimal surface for the shrinker metric) and connected. Thus (again by the compactness theorem [white-compact]*Theorem 1.1), the convergence of $\Sigma_{g}^{+}$ to $\Sigma^{+}$ is smooth and with multiplicity $1$, and therefore $K\cap\{z>0\}$ is empty. Therefore, by symmetry,
@@
Suppose $\Sigma^{+}$ is nonempty. Note that
(48) $\overline{\Sigma^{+}}\cap\{z=0\}\neq\emptyset,$
as otherwise $\Sigma^{+}$ would be a smooth shrinker (without boundary) in $\{z>0\}$. Note also that $\overline{\Sigma^{+}}$ is a surface of revolution. Let $C^{\prime}$ be a horizontal circle in $\Sigma^{+}$. Let $C_{g}^{\prime}$ be smooth, simple closed curves in $\Sigma_{g}$ that converge smoothly to $C^{\prime}$. By Lemma 60, $C_{g}^{\prime}$ bounds a region $E_{g}$ in $\Sigma_{g}$ that is either a disk or an annulus, that lies in $\{z>0\}$, and that is properly embedded in $\mathbf{R}^{3}$. Passing to the limit, we see that $C^{\prime}$ bounds a region $E$ in $\Sigma^{+}$ that is either a disk or an annulus, and that is properly embedded in $\mathbf{R}^{3}$.
By the strong maximum principle, $E$ lies in $\{z>0\}$. Since $E$ is properly embedded in $\mathbf{R}^{3}$,
(49) $\overline{E}\cap\{z=0\}=\emptyset.$
Now consider the other portion $E^{\prime}$ of $\Sigma^{+}$ bounded by $C^{\prime}$. By (48) and (49),
(50) $\overline{E^{\prime}}\cap\{z=0\}\neq\emptyset.$
Note that $E^{\prime}$ is a surface of revolution generated by a curve $e^{\prime}$ in the plane $y=0$, and that $e^{\prime}$ has one endpoint $p$ in $C^{\prime}\cap\{y=0\}$. By (50), the closure of $e^{\prime}$ contains a point $q$ in $\{z=0\}$. Since $\Sigma$ has finite entropy, the curve $e^{\prime}$ has locally finite length. Thus $e^{\prime}$ is a curve joining $p$ to $q$.
Now $q$ cannot be the origin, since otherwise $\Sigma^{+}\cup\{0\}$ and $\{z=0\}$ would violate the strong maximum principle at the origin. (One can invoke various versions of the maximum principle here. For example, the classical strong maximum principle suffices, because $\Sigma^{+}\cup\{0\}$ is a smooth minimal surface with respect to the shrinker metric by [structure]*Corollary 2.7.)
Thus $q=(x,0,0)$ for some $x\neq 0$. Therefore, $\overline{\Sigma^{+}}$ is a properly embedded disk or annulus whose boundary is the circle $C(r)$ (where $r=|x|$) in the plane $\{z=0\}$. This completes the proof of Claim 1. ∎
@@
Since $X\subset Q_{g}\subset\Sigma_{g}$, it follows that $p\in\Sigma_{g}$. By rotational symmetry by $\pi$ about $X$, if $L$ is a closed vertical line segment centered at $p$, then $L\cap\Sigma_{g}$ has an odd number of points (namely, the point $p$ and equal numbers of points above and below $p$.) ∎
@@
$m\mathbf{e}_{1}+m^{\prime}\mathbf{e}_{1}+\nu^{+}+\nu^{-}=0.$
@@
If $m=1$, the convergence is smooth by Allard’s Regularity Theorem. Thus we may assume that $m>1$.
For all sufficiently large $g$, $\Sigma_{g}\cap\partial B(0,r)$ consists of $m$ simple closed curves. Let $\Gamma_{g}^{0}$ be the curve that passes through the points of $Q_{g}\cap\partial B(0,r)$. Let $\Gamma_{g}$ be one of the other curves. By symmetry, we may assume that $\Gamma_{g}$ lies in $\{z>0\}$. Let $\Sigma_{g}^{\prime}$ be the component of $\Sigma_{g}\setminus\Gamma_{g}$ that lies in $\{z>0\}$. (The component exists by Lemma 60.) Since $\Sigma_{g}\setminus B(0,r)$ is connected and contains points in $\{z\leq 0\}$, we see that $\Sigma_{g}^{\prime}$ is contained in $B(0,r)$.